OpenStudy (anonymous):

diffrerence quotient f(x) =2xsquared + 4x -3

OpenStudy (anonymous):

\[\frac{2(x+h)^2+4(x+h)-3-2x^2-4x+3}{h}\] \[=\frac{2(x^2+2xh+h^2)+4x+4h-3-2x^2-4x+3}{h}\] \[=\frac{2x^2+4xh+2h^2+4x+4h-3-2x^2-4x+3}{h}\] \[=\frac{4xh+2h^2+4h}{h}=\frac{h(4x+2h+4)}{h}=4x+2h+4\]